[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Mar 20, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Mon Mar 20 15:41:44 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig

----------

TAFB:
1. TSB and NCO are continuing to upgrade the N-AWIPS systems to the RHEL 6  
operating system. The Sheephead backup workstation can be used at the  
Surface Analyst desk during this upgrade process.

2. The Cumulus workstation is down due to a failure of its CD-ROM drive.  
NCO Linux Sysadmin Brian Maher reports that it will be replaced within 24  
hours.

3. The Himawari imagery is available again in N-AWIPS. TAFB forecasters are  
encouraged to mention problems in their shift logs or let Monica Bozeman  
know directly if anything looks weird to assist TSB in monitoring this data.

4. Doug Gaer and the NCEP/EMC Wave Modeling Group are increasing the size  
of the NWPS grid domains to match the RWPS grid domains so more precise  
performance comparisons can be conducted between the two systems. AWIPS II  
Sysadmin Chris Mello will coordinate with TAFB later this week to get the  
new domains into GFE.

5. Windows Sysadmin Kim Gordon has removed the situational awareness system  
to address an intermittent fan problem (loud noise). She is setting up a  
backup system for TAFB that should be ready by COB today.
HSU:
1. AWIPS II Sysadmin Chris Mello will begin replacing the AWIPS printers  
and upgrading ten AWIPS monitors in HSU and in the Fish Bowl tomorrow,  
Tuesday, March 21.

2. The Jason-2 spacecraft entered safe mode on Wednesday, March 15th due to  
a single-event glitch in Gyrometer 1, immediately interrupting its  
measurements. Full mission recovery should not be expected before Thursday,  
March 23rd.
Other items:

NHC RFCs:
1. RFC #221 - dbn_send_daemon v1.3.0 - Add a subroutine to check for KML  
files in the "kml" directory. The Alert Type will be set to "DATA" and the  
Alert Name will be set to "NHC_KML". This uses the same methodology as  
existing file types. KML files contain storm surge watches/warnings. The  
files are sent to NCO. Implemented on Monday, March 20 at 1200Z (8 am EDT).

2. RFC #3033 - Make routing changes for NHC to add Virtual Routing and  
Forwarding (VRF) capability so the WFO-MIA and NHC networks may be  
combined. To be implemented on Wednesday, March 22 at 1300Z (9 am EDT).

3. RFC #3070 - Modified the parsing table on the production WCOSS to send  
Para data from the PSURGE model to NHC's DBN1/2 systems and DBN1awips  
system for testing purposes. This is being done in preparation for the  
PSURGE model upgrade in early April. Implemented on Thursday, March 16 at  
1730Z (1:30 pm EDT).

4. RFC #3020 - IBM will make a change to the 'access.conf' file on all  
access/login nodes on WCOSS to allow a new ATCF server (atcfsvr3) at NHC to  
access WCOSS, so NHC may run their guidance suite from that new server. To  
be implemented on the production WCOSS machine on Monday, March 20 at 1730Z  
(1:30 pm EDT).

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170320/fda2cadd/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list